OpenStudy (anonymous):

Given the proportion, a/b=2/17 , what ratio completes the equivalent proportion a/2=?/? Question 16 options: a) a/17 b) 17/b c) 2/17 d) b/17

OpenStudy (rsadhvika):

thats the complete question ?

OpenStudy (anonymous):

ok i fixed it

OpenStudy (rsadhvika):

\(\large \frac{a}{b}=\frac{2}{17} \) simply think of swapping the corners, 2 and b \(\large \frac{a}{2}=\frac{b}{17} \)

OpenStudy (anonymous):

ok thanks :DD

OpenStudy (anonymous):

can you help with one more question like this

OpenStudy (anonymous):

i guess its kinda like this

OpenStudy (rsadhvika):

think kinda questions you may try to solve like this, give it a try, if u get stuck il give a hand :)

OpenStudy (rsadhvika):

when u have a proportion u can swap corners

OpenStudy (rsadhvika):

lets do one more q, then u may understand better

OpenStudy (anonymous):

im always stuck on something

OpenStudy (anonymous):

ill post the last question

OpenStudy (rsadhvika):

its okay, i feel im stuck on my problems all the time...

OpenStudy (rsadhvika):

okay